Solve for the equation 5/2t= -3/5 I get 2=t is this true?
Answer by xcentaur(357)

5/2t= -3/5
multiply by 5
25/2t=-3
multiply both sides by 2t
25=-3*2t
25=-6t
t=-25/6


Note that the answer will change if the sign on the left hand side changes


Hope this helps,
Prabhat